OpenStudy (dallaslifebaby01):

What is the equation of the line in standard form that passes through the point (1, 24) and has a slope of -0.6? a) 3x + 5y = 125 b) 3x + 5y = 77 c) 3x + 5y = 123 d) 3x + 5y = 115

zepdrix (zepdrix):

How bout start by writing your line in `point-slope` form,\[\large\rm y-y_1=m(x-x_1)\]Plugging in your point,\[\large\rm y-24=m(x-1)\]And your slope,\[\large\rm y-24=-0.6(x-1)\]

zepdrix (zepdrix):

Try to shimmy some stuff around to get it in standard form :d

OpenStudy (dallaslifebaby01):

okay thanks! @zepdrix

OpenStudy (dallaslifebaby01):

standard form would be 0.6x+y=24.6

OpenStudy (dallaslifebaby01):

@zepdrix

zepdrix (zepdrix):

Mmm I guess I should have given you another hint. Fractions and decimals are not fun to work with. \(\large\rm y-24=-0.6(x-1)\) So as a first step, it might be a good idea to change the -0.6 to a whole number. We can achieve this by multiplying both sides by 10, right? That should move the decimal 1 place, giving us -6. \(\large\rm 10y-240=-6(x-1)\) Actually you already did a bunch of work, \(\large\rm 0.6x+y=24.6\) so let's just multiply by 10 from the point you have reached, \(\large\rm 6x+10y=246\) Confusing? What do you think? :d

OpenStudy (dallaslifebaby01):

so i would get c as my answer?

zepdrix (zepdrix):

Yayyy good job \c:/

OpenStudy (dallaslifebaby01):

thanks

zepdrix (zepdrix):

If you have some free time you should go clean those mandarin oranges out of your hair. They belong in a salad or ice cream, but not hair silly! Buhaha I'm just messin :p np

OpenStudy (dallaslifebaby01):

haha